OpenStudy (anonymous):

Segment TU has endpoints T(0 , -3) and U(0, 1). Select the equation in point-slope form for the perpendicular bisector of TU.

OpenStudy (anonymous):

A. y-1=2 B. y+1=0 C. y+1=-2 D. y-1=0

ganeshie8 (ganeshie8):